# STRAP: Robot Sub-Trajectory Retrieval for Augmented Policy Learning

### v1

@@ A.1 Simulation Experiments
BC $28.0\%\pm 0.9$ $0.0\%\pm 0$ 0.0 $17.3\%\pm 4.5$ $26.7\%\pm 4.3$ $18.0\%\pm 2.5$
Fine-Tuning $\bm{38.0\%\pm 5.7}$ $0.0\%\pm 0$ 0.0 $5.0\%\pm 2.1$ $\bm{81.0\%\pm 3.5}$ $\bm{35.0\%\pm 3.5}$
Multi-Task $10.0\%\pm 7.1$ $0.0\%\pm 0$ 0.0 $\underline{24.0\%\pm 17.0}$ $\underline{73.0\%\pm 13.4}$ $9.0\%\pm 2.1$
BR Du et al. (2023) $28.7\%\pm 3.9$ $0.0\%\pm 0$ 0.0 $13.3\%\pm 3.8$ $32.0\%\pm 4.3$ $26.0\%\pm 1.9$
FR Lin et al. (2024) $27.3\%\pm 1.4$ $0.0\%\pm 0$ 0.0 $11.3\%\pm 3.0$ $41.3\%\pm 5.5$ $14.7\%\pm 1.1$
D-S $30.0\%\pm 3.4$ $0.0\%\pm 0$ 0.0 $4.7\%\pm 0.5$ $16.0\%\pm 5.7$ $6.0\%\pm 0.9$
D-T $\underline{34.7\%\pm 2.0}$ $0.0\%\pm 0$ 0.0 $4.7\%\pm 1.1$ $27.3\%\pm 4.5$ $14.0\%\pm 3.4$
STRAP (CLIP, $K{=}100$) $30.0\%\pm 2.5$ $0.0\%\pm 0$ 0.0 $8.7\%\pm 6.3$ $29.3\%\pm 10.5$ $24.0\%\pm 4.3$
STRAP (DINOv2, $K{=}100$) $29.3\%\pm 2.7$ $0.0\%\pm 0$ 0.0 $16.7\%\pm 2.0$ $29.3\%\pm 11.3$ $18.7\%\pm 1.4$
STRAP (DINOv2, best $K$) $32.0\%\pm 5.7$ $0.0\%\pm 0$ 0.0 $\bm{61.0\%\pm 6.4}$ $61.0\%\pm 0.7$ $\underline{31.0\%\pm 2.1}$
@@
Following  Lin et al. (2024), we retrieve single-state action pairs for the state-based retrieval baselines (BR, FR, D-S) and pad them by also retrieving the states from $t-h$ to $t+h-1$ to make the samples compatible with our transformer-based policy. We refer the reader to Sec. A.3 for extensive ablation.
@@ A.3 Ablations
Sub-traj $76.0\%\pm 4.71$ $75.33\%\pm 2.72$ $26.0\%\pm 1.89$ $00.0\%\pm 0$ 0.00 $37.33\%\pm 6.62$
Full traj $78.67\%\pm 2.72$ $68.67\%\pm 1.44$ $34.67\%\pm 1.96$ $00.0\%\pm 0$ 0.00 $28.67\%\pm 3.81$

### v2

@@ A.1 Simulation Experiments
BC $28.0\%\pm 0.9$ $0.0\%\pm 0.0$ $17.3\%\pm 4.5$ $26.7\%\pm 4.3$ $18.0\%\pm 2.5$
Fine-Tuning $\bm{38.0\%\pm 5.7}$ $0.0\%\pm 0.0$ $5.0\%\pm 2.1$ $\bm{81.0\%\pm 3.5}$ $\bm{35.0\%\pm 3.5}$
Multi-Task $10.0\%\pm 7.1$ $0.0\%\pm 0.0$ $\underline{24.0\%\pm 17.0}$ $\underline{73.0\%\pm 13.4}$ $9.0\%\pm 2.1$
BR Du et al. (2023) $28.7\%\pm 3.9$ $0.0\%\pm 0.0$ $13.3\%\pm 3.8$ $32.0\%\pm 4.3$ $26.0\%\pm 1.9$
FR Lin et al. (2024) $27.3\%\pm 1.4$ $0.0\%\pm 0.0$ $11.3\%\pm 3.0$ $41.3\%\pm 5.5$ $14.7\%\pm 1.1$
D-S $30.0\%\pm 3.4$ $0.0\%\pm 0.0$ $4.7\%\pm 0.5$ $16.0\%\pm 5.7$ $6.0\%\pm 0.9$
D-T $\underline{34.7\%\pm 2.0}$ $0.0\%\pm 0.0$ $4.7\%\pm 1.1$ $27.3\%\pm 4.5$ $14.0\%\pm 3.4$
STRAP (CLIP, K=100K{=}100) $30.0\%\pm 2.5$ $0.0\%\pm 0.0$ $8.7\%\pm 6.3$ $29.3\%\pm 10.5$ $24.0\%\pm 4.3$
STRAP (DINOv2, K=100K{=}100) $29.3\%\pm 2.7$ $0.0\%\pm 0.0$ $16.7\%\pm 2.0$ $29.3\%\pm 11.3$ $18.7\%\pm 1.4$
STRAP (DINOv2, best $K$) $32.0\%\pm 5.7$ $0.0\%\pm 0.0$ $\bm{61.0\%\pm 6.4}$ $61.0\%\pm 0.7$ $\underline{31.0\%\pm 2.1}$
@@
Following  Lin et al. (2024), we retrieve single-state action pairs for the state-based retrieval baselines (BR, FR, D-S) and pad them by also retrieving the states from tht-h to t+h1t+h-1 to make the samples compatible with our transformer-based policy. We refer the reader to Sec. A.3 for extensive ablation.
@@ A.3 Ablations
Sub-traj $76.0\%\pm 4.71$ $75.33\%\pm 2.72$ $26.0\%\pm 1.89$ $0.0\%\pm 0.00$ $37.33\%\pm 6.62$
Full traj $78.67\%\pm 2.72$ $68.67\%\pm 1.44$ $34.67\%\pm 1.96$ $0.0\%\pm 0.00$ $28.67\%\pm 3.81$